The dimesions of a rectangular fish pond are 6m and 9m. The pond is surrounded by a cement walk of uniform width. If the area of the walk is equal to the area of the fish pond,write an equation in the simplified form showing this information.
Answer by mananth(16949)

let the width be x
length of pond plus width of walk way = 9+2x
width of pond + walkway = 6+2x
Area of (pond + walkway)= (9+2x)(6+2x)
(9+2x)(6+2x)-54 = 54
simplify to get the solution